Fill in the gaps in the table below

Quantity of           Total Output     Marginal Product     Average product of
Variable of input                      Variable input        Variable input
0                        0                   -                  -
1                        225
2                                                               300
3                                           300
4                        1140
5                                           225
6                                                                225

    
1
Expert's answer
2021-12-24T07:31:40-0500

Working notes:

Total output of input 2 = Average product × 2 = 300 × 2 = 600

Total output of input 6 = Average product × 6 = 225 × 6 = 1350

Marginal product input 1 = Total output of input 1 – Total output 0 = 225 - 0 = 225

Marginal product input 2 = Total output of input 2 – Total output 1 = 600 - 225 = 375

Marginal product input 4 = Total output of input 4 – Total output 3 = 1140 - 900 = 240

Marginal product input 6 = Total output of input 6 – Total output 5 = 1350 - 1365 = -15